OpenStudy (marsxtc):

Calculus Derivatives : a projectile is fired vertically upward with an initial velocity of 16 ft/sec. 1)How long will it take for the projectile to hit the ground? 2) how high will it go? I thought it would use the equation S(0)+V(0)t- (1/2)gt^2 but I'm not sure what to do at the moment. S(0) Initial position V(0) initial velocity and g = gravity

OpenStudy (phi):

S(t)= S(0)+V(0)t- (1/2)gt^2 put in the numbers \[ S(t)= 0 + 16 t - \frac{1}{2} 32 t^2 \\ S(t)= - 16 t^2+16 t \] this is a parabola, so you can find its vertex (t= -b/(2a) ) or use calculus to find where the slope is 0 you can find the total flight time by solving for S(t)=0 or you can double the time it takes to reach the peak (it takes as much time to go up as down)

OpenStudy (anonymous):

Phi, I thought you had it right the first time. Can you explain to me where the 32 came from for g?

OpenStudy (marsxtc):

Since S(T)= -16t^2+16t that counts as its position. and then Velocity would be = -32t+16 Acceleration would be -32. Wouldn't setting the velocity to 0, would give me the time as when its at its highest peak, showing me the largest height? 0=-32t+16 = t= -2. And then putting it into S(T) -16(-2)^2 + 16(-2) Would lead me to when its still at its highest point showing me how high it will go?

OpenStudy (marsxtc):

@jb1515g gravity I believe is 32.174 ft/s2 so thats where he got 32

OpenStudy (anonymous):

Ah, I see now. I didn't notice it was in feet. My mistake. I was working in meters.

OpenStudy (phi):

yes, I used meters/sec^2 for g in the first post, and then noticed the problem is using feet for units

OpenStudy (anonymous):

Yeah, I did the same thing, just didn't notice it.

OpenStudy (marsxtc):

Would setting S(T) = 0 = -16t^2+16t equal to how long it will take for the projectile to hit the ground?

OpenStudy (anonymous):

Check your math, marsxtc. In 0=-32t+16, t isn't -2.

OpenStudy (marsxtc):

Ah I see, I messed up a bit I wast trying to do it fast. So it'd be 1/2.

OpenStudy (anonymous):

No, taking S(T) = -16t^2 + 16t and plugging in t will give you the max height. To find how long it would take to hit the ground, you take the t you just found (which is the time it takes to get to the max height) and double it, since it will take the same time to fall from that height. *edit for clarification

OpenStudy (phi):

*** Would setting S(T) = 0 = -16t^2+16t equal to how long it will take for the projectile to hit the ground? *** or solving for t when the height is 0 (i.e. on the ground) we find \[ t (-16t + 16) = 0 \\ t= 0 \text{ or } -16t+16=0 \\ -16t= -16 \\ t = 1 \] so t=0 (at launch) and t= 1 (back on the ground)

OpenStudy (marsxtc):

So the time is 1/2 for it to reach maximum, so 1 second for it to reach the ground?

OpenStudy (anonymous):

Yes sir.

OpenStudy (marsxtc):

So according to phi plugging in 0 as the position would also work wouldn't it?

OpenStudy (anonymous):

Yeah, that works too. I always found finding for max height, then doubling, easier, But both work.

OpenStudy (marsxtc):

Ok just to clarify ( I have a huge test on derivatives and I forgot how to do everything related to this if theres no like equation given to me) Finding the S(T), I can determine the V(T) and the A(T). Setting V(T) to 0 Will give me the time for the maximum height because at velocity = 0 its yeah isn't moving. inputting that time into S(T) will give me the height? And setting S(T) to 0 will give me the time of when it hits the ground?

OpenStudy (anonymous):

Yes, that is correct.

OpenStudy (marsxtc):

Alright. I really don't know on who to give the medal to. Both of you were so helpful!...

OpenStudy (anonymous):

I'd say Phi is deserving

OpenStudy (marsxtc):

Alright thank you for taking your time to help me!